OpenStudy (anonymous):

A company distributes free candies to all students of x schools. Each school has (x+1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A) Write an expression to show the total number of candies distributed by the company in x schools. Part B) What would x(x+1) represent? When simplified, what would be the degree and classification of this expression? Part C) How can you calculate the total number of students in each school? I only need help for Part C.

OpenStudy (anonymous):

My answers: part a) the total number of candies is given 4 each to x+4 students in x+1 classes in x schools. So total number of candies = 4*(x+4)*(x+1)*x part b)x is the number schools x+1 is the number of classes in each school x*(x+1) would represent the total number of classes all taken together. When simplified x(x+1) = x2+1 which is a second degree polynomial.

OpenStudy (anonymous):

@phi @iGreen @dan815

OpenStudy (phi):

I am thinking they want the expression for total number of students in each school which is (x+1) classes/school * (x+4) students/class = (x+1)(x+4) students/school so (x+1)(x+4) students/school